OpenStudy (anonymous):

Surf's Up charges $22 per hour to rent a surfboard and $8 per hour to rent a wetsuit. Marty and Cheryl paid a total of $240 to rent a surfboard and a wetsuit each for a few hours one day. Cheryl rented the surfboard and wetsuit for 2 hours less than Marty. The following equation represents this situation, where x represents the number of hours Marty rented the surfboard and the wetsuit. 22x + 8x + 22(x − 2) + 8(x − 2) = 240 What is the total amount Marty paid to rent the surfboard and wetsuit that day?

OpenStudy (mertsj):

30x+22x-44+8x-16=240 60x-60=240 60x=300 x=5 So Marty rented the surfboard and wetsuit for 5 hours and paid 22(5)=8(5)=$150